# AttentionX: Exploiting Consensus Discrepancy In Attention from A Distributed Optimization Perspective

### v1

@@ 1 Introduction
In recent years, transformers [15] in artificial intelligence have achieved great success in various data-analysis domain such as natural language processing (NLP) [2; 14], computer vision [6], image generation and editing [11; 7; 19], and audio processing [9]. One key component in transformer is the Attention layer for capturing long-distance dependency across a sequence of tokens. Specifically, the Attention operation gathers relevant information from other tokens for each one via the $KQ$-softmax-based weighted summation of $V$-representations. The feedforward network (FFN) further processes the output of the Attention layer per token, which can be interpreted as local information fusion.
@@
In this paper, we aim to extend the standard Attention by following a similar procedure as PDMM. In particular, we compute the consensus discrepancy in Attention as the difference between the weighted summation of the $V$-representations and the scaled $V$-representations themselves. It is hypothesized that as the index of the Attention-FFN layer increases, the computed consensus discrepancy tends to decrease and become stable as the dependency across all the tokens would be fully captured when the transformer gets very deep. This is conceptually similar to the consensus achieved by PDMM to a certain extent, of which the residual errors of the linear edge-constraints gradually decreases as the iteration of PDMM increases. Similarly to PDMM, we propose to incorporate the above consensus discrepancy in computing the output of Attention, referred to as AttentionX. Experiments indicate that AttentionX indeed improves the validation performance of ViT and nanoGPT.
@@ 2 Brief Review of PDMM
Detailed convergence results of the algorithm can be found in [18; 12].
@@ 3.1 Revisiting Attention-FFN framework in transformer
The original work [15] proposes the encoder-decoder structure in the transformer for NLP applications. The Attention-FFN framework is slightly different in encoder and decoder. For demonstration purpose, we consider a simplified version, represented as (see [1; 6])
$\textrm{head}_{i}(X)$ $=\textrm{Attention}(\overbrace{XW_{i}^{Q}}^{Q_{i}},\overbrace{XW_{i}^{K}}^{K_{i}},\overbrace{XW_{i}^{V}}^{V_{i}})$ (7)
@@
$X$ $\leftarrow\underbrace{\textrm{FFN}(\textrm{MultiHead}(X)+X)+\textrm{MultiHead}(X)+X}_{\textrm{{\color[rgb]{0,0,1}info. fusion}}},$ (9)
where $(W_{i}^{Q},W_{i}^{K},W_{i}^{V})$ are the three learnable matrices for computing $(Q_{i},K_{i},V_{i})$ of the $i$th attention, and Concat stacks up $h$ attentions $\{\textrm{head}_{i}(X)\}_{i=1}^{h}$, which is multiplied by the learnable matrix $W^{o}$ as the output of the Attention layer.
@@
$\textrm{head}_{i}(X)=\overbrace{\textrm{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{i}}}\right)V_{i}}^{\textrm{{\color[rgb]{0,0,1}info. gathering}}},$ (10)
where $d_{i}$ is the dimension of the $Q_{i}$ vectors. The softmax term computes the unified relevance of each token with respect to neighbouring tokens. Similarly to that of PDMM, the computed weighted summation of $V$ representations can be taken as information gathering from all neighbours.
@@ 3.2 Update expression of AttentionX
As reviewed earlier, PDMM exploited consensus discrepancy in the form of residual errors of the linear equality constraints in its update expressions. Similarly, we extend Attention by also incorporating the associated consensus discrepancy. We let the consensus discrepancy in the context of Attention as the difference between the weighted summation of $V$ representations and scaled $V$ representations themselves, given by
$\Phi_{i}(X)=\gamma V_{i}-\textrm{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{i}}}\right)V_{i}\quad i=1,\ldots,h,$ (11)
where $\gamma\in(0,1]$. We hypothesize that as the transformer goes deeper, the dependency across the tokens would be gradually captured by the softmax operation. As a result, the magnitude of consensus discrepancy $\Phi_{i}$ would decrease and become stable eventually. Inspired by PDMM, we compute the output of the multi-head attention in terms of $\{\Phi_{i}\}_{i=1}^{h}$ as
@@
We refer to (7) and (9), together with the new update expressions (11)-(12) as AttentionX to differentiate it from the classical Attention. It is clear that no additional learnable parameters are introduced in AttentionX. The computational overhead in (11) only involves one subtration operation, which is negligible.
Finally, we briefly explain why the consensus discrepancy from earlier AttentionX layers are not utilized for the computing the output of current AttentionX layer. We first note that in PDMM, the Lagrangian multipliers $\{\lambda_{j|i}\}_{j\in\mathcal{N}_{i}}$ accumulate all the associated historical residual errors, which are then used for updating the primal variable $\boldsymbol{x}_{i}$. Intuitively speaking, this is because the linear edge-constraints are fixed over iterations. The historical residual errors play an role in the current iteration. On the other hand, the dependency across tokens in transformers is dynamically learned by stacking a set of Attention-FFN layers. There is no explicit and fixed constraint between tokens. Therefore, in AttentionX, we only employ the most recent consensus discrepancy as specified in (11)-(12).
@@ 4.1 On training ViT-small
In this experiment, we consider training ViT-small over CIFAR10 and CIFAR100. The $\gamma$ parameter in AttentionX (see (11)) was set to $\gamma=1$. The SET-Adam optimizer was utilized [17] in the training process with the configuration $(\eta_{0},\beta_{1},\beta_{2},\epsilon)=(1e-4,0.9,0.999,1e^{-18})$, where $\eta_{0}$ denotes the initial learning rate. The remaining training setups follow directly from the original open source. Three experimental repetitions were performed per training setup to mitigate the effect of randomness.
Table 1 summarizes the obtained validation accuracy. It is clear that ViT-small with AttentionX produces considerably better performance than with Attention. This indicates that the consensus discrepancy characterised by $\Phi_{i}(X)$ in (11) is a better choice than the softmax-based $V$-representations.
@@ 4.2 On training nano-GPT2
In this experiment, we consider training nano-GPT2 by using the dataset of openwebtext. To save training time, we only took a small subset from the entire training dataset when training the model. The parameter $\gamma$ in attentionX was set to $\gamma=0.5$.
@@ 5 Conclusions
In this work, we have proposed AttentionX to replace Attention in transformer from a distributed optimization perspective. In particular, we first identify similarity between the update expressions of PDMM and the classical Attention-FFN framework in transformer. The Attention operation can be viewed as information gathering (corresponding to message passing in PDMM) from neighbouring tokens while the FFN operation can be taken as local information fusion (corresponding to the node-oriented optimization in PDMM). Inspired by PDMM that exploits the consensus discrepancy in its update expressions, we also utilize the consensus discrepancy in the form of the difference between the weighted summation of $V$-representations and the scaled $V$-representations themselves when designing AttentionX. One additional learnable parameters are needed AttentionX. Experiments on ViT-small and nano-GPT2 show that AttentionX leads to better validation performance.

### v2

@@ 1 Introduction
In recent years, transformers [24] in artificial intelligence have achieved great success in various data-analysis domain such as natural language processing (NLP) [3; 23], computer vision [7], image generation and editing [20; 8; 29], and audio processing [13]. One key component in transformer is the Attention layer for capturing long-distance dependency across a sequence of tokens. Specifically, the Attention operation gathers relevant information from other tokens for each one via the $KQ$-softmax-based weighted summation of $V$-representations. The feedforward network (FFN) further processes the output of the Attention layer per token, which can be interpreted as local information fusion.
@@
In this paper, we aim to extend the standard Attention by following a similar procedure as PDMM. In particular, we compute the consensus discrepancy in Attention as the difference between the scaled weighted summation of the $V$-representations and the $V$-representations themselves. It is hypothesized that as the index of the Attention-FFN layer increases, the computed consensus discrepancy tends to decrease and become stable as the dependency across all the tokens would be fully captured when the transformer gets very deep. This is conceptually similar to the consensus achieved by PDMM to a certain extent, of which the residual errors of the linear edge-constraints gradually decreases as the iteration of PDMM increases. Similarly to PDMM, we propose to incorporate the above consensus discrepancy in computing the output of Attention, referred to as AttentionX. Experiments indicate that AttentionX indeed improves the validation performance of ViT and nanoGPT.
@@ 2 Brief Review of PDMM
Detailed convergence results of the algorithm can be found in [27; 21].
@@
In addition to the Lagrangian multipliers for capturing the historical consensus discrepancy, it is clear from (4) that the set of quadratic penalty functions $\{\|A_{ij}x_{i}+A_{ji}x_{j}^{k}-b_{ij}\|\}_{j\in\mathcal{N}_{i}}$ are also included in computation of $x_{i}^{k+1}$. The penalty functions attempt to softly constrain $x_{i}^{k+1}$ in a region that incurs small consensus discrepancy (with regard to the predefined edge-constraints) with respect to the neighbors $\{x_{j}^{k}\}_{j\in\mathcal{N}_{i}}$. The parameter $\rho>0$ in front of the penalty functions and in (6) controls the contribution of the consensus discrepancy when updating the primal variables $\{x_{i}\}_{i\in\mathcal{V}}$.
Since the invention of PDMM, it has received considered research investigation in the past few years. The work [10] studied the convergence of stochastic PDMM which includes asynchronous PDMM and PDMM with transmission losses between neighbours as special cases. In [28], PDMM is modified for federated learning over a centralised network, where it is found that PDMM is closely related to the SCAFFOLD [11] and FedSplit [19] algorithm. Additionally, PDMM can be employed for privacy-preserving distributed optimisation, providing a level of privacy assurance, by utilising the fact that the (synchronous) PDMM updates take place within a particular subspace and the orthogonal complement can be used to obscure local (private) data, a technique known as subspace perturbation [16; 14; 17; 15]. Additionally, research in [2] demonstrates that PDMM exhibits robustness against data quantisation. Recently, the PDMM algorithm has been extended to incorporate affine inequality constraints as well [9]. This enhancement enables its application in solving linear programs in a distributed fashion.
@@ 3.1 Revisiting Attention-FFN framework in transformer
The original work [24] proposes the encoder-decoder structure in the transformer for NLP applications. The Attention-FFN framework is slightly different in encoder and decoder. For demonstration purpose, we consider a simplified version, represented as (see [1; 7])
$\textrm{head}_{m}(X)$ $=\textrm{Attention}(\overbrace{XW_{m}^{Q}}^{Q_{m}},\overbrace{XW_{m}^{K}}^{K_{m}},\overbrace{XW_{m}^{V}}^{V_{m}})$ (7)
@@
$Y$ $=X+\textrm{MultiHead}(X)$ (9)
$X$ $=\underbrace{\textrm{FFN}(Y)+Y}_{\textrm{{\color[rgb]{0,0,1}info.fusion}}},$ (10)
where $(W_{m}^{Q},W_{m}^{K},W_{m}^{V})$ are the three learnable matrices for computing $(Q_{m},K_{m},V_{m})$ of the $m$th attention, and Concat stacks up $h$ attentions $\{\textrm{head}_{m}(X)\}_{m=1}^{h}$, which is multiplied by the learnable matrix $W^{o}$ as the output of the Attention layer.
@@
$\textrm{head}_{m}(X)=\overbrace{\textrm{softmax}\left(\frac{Q_{m}K_{m}^{T}}{\sqrt{d_{m}}}\right)V_{m}}^{\textrm{{\color[rgb]{0,0,1}info. gathering}}},$ (11)
where $d_{m}$ is the dimension of the $Q_{m}$ vectors. The softmax term computes the unified relevance of each token with respect to neighbouring tokens. Similarly to that of PDMM, the computed weighted summation of $V$ representations can be taken as information gathering from all neighbours.
@@ 3.2 Update expression of AttentionX
As reviewed earlier, PDMM exploited consensus discrepancy in the form of residual errors of the linear equality constraints in its update expressions. Similarly, we extend Attention by also incorporating the associated consensus discrepancy. We let the consensus discrepancy in the context of Attention as the difference between the scaled weighted summation of $V$ representations and $V$ representations themselves, given by
$\Phi_{m}(X)=V_{m}-\gamma\textrm{softmax}\left(\frac{Q_{m}K_{m}^{T}}{\sqrt{d_{m}}}\right)V_{m}\quad m=1,\ldots,h,$ (12)
where $\gamma\geq 1$. Because of the softmax-based weighted summation, more relevant tokens contribute more to the measured consensus discrepancy. Inspired by PDMM, we compute the output of the multi-head attention in terms of $\{\Phi_{m}\}_{m=1}^{h}$ as
@@
We refer to (7) and (9)-(10), together with the new update expressions (12)-(13) as AttentionX to differentiate it from the classical Attention. It is clear that no additional learnable parameters are introduced in AttentionX. The computational overhead in (12) only involves one subtration operation, which is negligible.
Finally, we briefly explain why the consensus discrepancy from earlier AttentionX layers are not utilized for the computing the output of the current AttentionX layer. We first note that in PDMM, the Lagrangian multipliers $\{\lambda_{j|i}\}_{j\in\mathcal{N}_{i}}$ accumulate all the associated historical residual errors, which are then used for updating the primal variable $\boldsymbol{x}_{i}$. Intuitively speaking, this is because the linear edge-constraints are fixed over iterations. The historical residual errors play an role in the current iteration. On the other hand, the dependency across tokens in transformers is dynamically learned by stacking a set of Attention-FFN layers. There is no explicit and fixed constraint between tokens. Therefore, in AttentionX, we only employ the most recent consensus discrepancy as specified in (12)-(13).
Regarding selection of $\gamma$ parameter in (12): We argue that the selection of the $\gamma$ parameter should depend on if the diagonal elements of the weighting matrix $\textrm{softmax}(Q_{m}K_{m}^{T}/\sqrt{d_{m}})$ are set to be zero or not. To simplify notations below, we use $\alpha_{k,j}$ to denote the $(k,j)$th element of the weighting matrix. Consider the $k$th row of $\Phi_{m}(X)$ for the $k$th token, which can be re-parameterized as
$\Phi_{m}(X)[k,:]$ $=V_{m}[k,:]-\gamma\sum_{j=1}^{n}\alpha_{k,j}V_{m}[j,:]$
$=\sum_{j=1}^{n}\alpha_{k,j}(V_{m}[k,:]-\gamma V_{m}[j,:])$
$=\underbrace{\alpha_{k,k}(1-\gamma)V_{m}[k,:]}_{{\color[rgb]{0,0,1}\textrm{1st term}}}+\underbrace{\sum_{j\neq k}\alpha_{k,j}(V_{m}[k,:]-\gamma V_{m}[j,:])}_{{\color[rgb]{0,0,1}\textrm{ 2nd term}}}$ (14)
where we assume there are $n$ tokens in total. In practice, one has the freedom to set $\alpha_{k,k}$ to be zero or not via the masking technique. By using (14) and the fact that $\sum_{j=1}^{n}\alpha_{k,j}=1$, one can easily show that when $\gamma=1$, there is
$\|\underbrace{\Phi_{m}(X)[k,:]}_{\gamma=1,\alpha_{k,k}=0}\|^{2}>\|\underbrace{\Phi_{m}(X)[k,:]}_{\gamma=1,\alpha_{k,k}\neq 0}\|^{2}.$ (15)
That is, the magnitude of the consensus discrepancy $\Phi_{m}(X)[k,:]$ decreases when $\alpha_{k,k}$ becomes non-zero without any masking.
To mitigate the magnitude difference in (15), we recommend to set $\gamma>1$ (or $\gamma=1$) when $\alpha_{k,k}\neq 0$ (or $\alpha_{k,k}=0$). The common practice is that in LLMs (e.g., nano-GPT), the diagonal elements are set to be nonzero by including the contributions from their own representations via the casual-masking.
@@ 4.1 On training ViT-small
In this experiment, we consider training ViT-small over CIFAR10 and CIFAR100. The $\gamma$ parameter in AttentionX (see (12)) was set to $\gamma=1$ as the diagonal elements of $\textrm{softmax}(Q_{m}K_{m}^{T}/\sqrt{d_{m}})$ are set to zero on purpose. The SET-Adam optimizer was utilized [26] in the training process with the configuration $(\eta_{0},\beta_{1},\beta_{2},\epsilon)=(1e-4,0.9,0.999,1e^{-18})$, where $\eta_{0}$ denotes the initial learning rate. The remaining training setups follow directly from the original open source. Three experimental repetitions were performed per training setup to mitigate the effect of randomness.
Table 1 summarizes the obtained validation accuracy. It is clear that ViT-small with AttentionX produces considerably better performance than with Attention. This indicates that the consensus discrepancy characterised by $\Phi_{m}(X)$ in (12) is a better choice than the softmax-based $V$-representations.
@@ 4.2 On training nano-GPT2
In this experiment, we consider training nano-GPT2 by using the dataset of openwebtext. To save training time, we only took a small subset from the entire training dataset when training the model. The parameter $\gamma$ in attentionX was set to $\gamma=3$.
@@ 5 Conclusions
In this work, we have proposed AttentionX to replace Attention in transformer from a distributed optimization perspective. In particular, we first identify similarity between the update expressions of PDMM and the classical Attention-FFN framework in transformer. The Attention operation can be viewed as information gathering (corresponding to message passing in PDMM) from neighbouring tokens while the FFN operation can be taken as local information fusion (corresponding to the node-oriented optimization in PDMM). Inspired by PDMM that exploits the consensus discrepancy in its update expressions, we also utilize the consensus discrepancy in the form of the difference between the scaled weighted summation of $V$-representations and the $V$-representations themselves when designing AttentionX. One additional learnable parameters are needed AttentionX. Experiments on ViT-small and nano-GPT2 show that AttentionX leads to better validation performance.